OpenStudy (anonymous):

Josh had $10 more than Carly, so he gave Carly half of his money. Then Carly had more than Josh, so she gave Josh $4. Then they both had $13.00. How much money did each person start with. Please show work.

OpenStudy (btaylor):

If both have 13 at the end, then there is $26 total. so write a system: J = 10 + C J + C = 26 substitute in J: 10 + C + C = 26 2C = 16 C = 8 so Carly had $8 originally. Josh had $10 more, so he had $18 originally.